# Subtleties in the pseudomodes formalism

### 04 Mar 2026 v2

@@ I Introduction
The success of the pseudomodes approach, however, relies on the careful design of their geometry and parameters so as to properly mimic the real environment. Implementing this entails various subtleties, which are often overlooked. The goal of this paper is to address said subtleties. In section II, we consider the effective spectral density that arises from a particular configuration of pseudomodes. Depending on the non-Hermitian Hamiltonian generated by the pseudomodes’ Hamiltonian and the residual baths, three possible cases emerge. If none of the pseudomodes couple to each other, then one obtains an effective spectral density that is a sum of Lorentzians, while if the pseudomodes couple to each other there are additional terms in the effective spectral density which we call “Anti-Lorentzian”. We take this one step further, and show that one can obtain terms which are neither Lorentzian nor Anti-Lorentzian by using a non-diagonalizable non-Hermitian Hamiltonian for the pseudomodes, a case which to the best of our knowledge has not been previously explored.
@@ I.1 Model System
$-\sum_{\alpha}\sum_{j\in S}\intop_{0}^{t}dt^{\prime}\chi_{\alpha,ij}(t-t^{\prime})c_{j}(t^{\prime})$ (5)
@@
$C^{+}_{\alpha,ij}(t)=\intop\frac{d\omega}{2\pi}e^{i\omega t}J_{\alpha,ij}(\omega)f_{\alpha}(\omega).$ (7)
$C^{-}_{\alpha,ij}(t)=\intop\frac{d\omega}{2\pi}e^{i\omega t}J_{\alpha,ij}(\omega)\big(1-f_{\alpha}(\omega)\big)\,$, (8)
@@
$\sum_{\alpha}\sum_{i\in S}c_{i}^{\dagger}B_{\alpha i}+B_{\alpha i}^{\dagger}c\,$, (9)
@@ I.2 Pseudomode Mapping
The first line of (I.2) describes the possible interactions between pseudomodes of the same bath, described by the $n_{\alpha}\times n_{\alpha}$ matrix $\Lambda_{\alpha}$ for each bath. If $\Lambda$ is diagonal then none of the pseudomodes couple to each other, and we have the simpler configuration shown in Fig. 1(b). The second line describes how each pseudomode interacts with the central system through a set of $n_{\alpha}\times n_{S}$ matrices $\zeta_{\alpha}$. Finally, the last line describes a set of infinite dimensional residual reservoirs, one for each pseudomode $a_{\alpha k}$, labeled by the pair of indices $(\alpha,k)$ and composed of a set of modes $b_{\alpha k,p}$. The residual baths then have spectral densities given by
@@ II Computation of Effective Spectral Density
While all the results shown in this section follow straightforwardly from (15) and (17), we show them explicitly, as we will use these results for the remainder of the paper. Writing these results explicitly also makes clear the difference in the effective spectral densities that can be obtained using coupled modes as opposed to uncoupled modes, and why it is often desirable to use coupled modes. Additionally, the case where $W_{\alpha}$ is not diagonalizable has not, to the best of our knowledge, been previously explored.
@@ II.2 Diagonalizable Case
Using a diagonalizable but not diagonal $W$ often allows for a much better fit to structured spectral densities than using a diagonal $W$, as the negativities and anti-symmetry from the Anti-Lorentzian terms allow one to capture features much more easily. Fig. 2 compares 6-mode fits to the semi-elliptical spectral density, showing how the diagonalizable version matches the true spectral density much more closely, particularly around the sharp corners at $\omega=\pm 1$.
@@ II.3 Non-Diagonalizable Case
Example: 2x2 non-diagonalizable Case
Consider the case where we map bath $\alpha$, which couples to only one system site (so the spectral density matrix is 1x1) to two pseudomodes, so $W_{\alpha}$ is 2x2. If we want $W_{\alpha}$ to be non-diagonalizable, the most general form we can pick is:
@@ III Few-Mode Fitting
The most common approach to few-mode fitting is to first fit the memory kernel in terms of the parameters $\kappa_{\alpha k,ij},\varepsilon_{\alpha k},\gamma_{\alpha k}$, and then to try to choose pseudomode parameters $\Lambda,\Gamma,\zeta$ that give rise to the same effective memory kernel. We focus on this method. One might alternately choose to fit the correlation functions (7) (or their Fourier transforms), thus encoding the temperatures and chemical potentials of the original baths into the pseudomode parameters, rather than the residual baths Pleasance et al. (2020); Menczel et al. (2024). Few-mode fitting has also been performed by brute- force optimization over the pseudomode parameters without first fitting the memory kernel Feist et al. (2020); Medina et al. (2021); Sánchez-Barquilla et al. (2022); doing so guarantees a fit with physical pseudomode parameters (which we shall see is not always guaranteed otherwise), but is much less efficient computationally. A third method was recently proposed in Ref. Huang et al. (2025), which instead maps (17) onto the so-called control-theoretic realization problem Mayo and Antoulas (2007); Antoulas and Anderson (1986).
@@ III.1 Fitting the Memory Kernel
We fit the memory kernel from each bath separately. We first fit each element of $\chi_{\alpha,ij}^{\rm eff}(t)$ as a sum of complex exponentials using techniques from discrete signal processing, such as Prony’s method de Prony (1795); Wilson et al. (2019); Almunif et al. (2020); Chen et al. (2022) or the ESPRIT algorithm Roy et al. (1986); Paulraj et al. (1986); Sahnoun et al. (2017) (see Appendix B). These methods work by sampling the true $\chi(t)$ at discrete, evenly spaced times, and then fitting those discrete points with complex exponentials. We found the best results by using ESPRIT, as it is more stable with respect to the set of points at which $\chi$ is sampled than Prony’s method. This gives us the set of $\kappa_{\alpha k,ij},\varepsilon_{\alpha k},\gamma_{\alpha k}$ from which to construct the effective memory kernel (24). Since the exponents in $\chi_{\alpha,ij}^{\rm eff}$ are independent of $i,j$, in the end each $i,j$ element of the memory kernel matrix should be fit by a sum of exponentials with the same exponents but different amplitudes $\kappa_{\alpha k,ij}$. Caution must be used here, since the fit obtained may produce a spectral density that is not positive everywhere or even has non-decaying exponentials. Note, also, that some spectral densities will be easier to fit than others. In particular, if we have a wideband $J(\omega)=\Gamma_{0}$ for $\omega_{m}\leq\omega\leq\omega_{M}$ and 0 otherwise, then the corresponding memory kernel is
@@ III.2 Obtaining Pseudomode Parameters
where $M$ is a diagonal matrix whose entries are the eigenvalues of $W$ obtained from the fitting the memory kernel. While the eigenvalues of $W$ are specified, $S$ is not. Note that if we transform $W$ by some unitary transformation, then as long as we transform $\zeta$ by the same unitary we will obtain the same effective memory kernel (24):
@@
It is worth emphasizing again here the sheer amount of freedom involved in this process. In (44) we can choose anything for the $u_{j}$. Then in (45), $B_{\perp}$ can by any positive-definite $n\times n$ matrix satisfying $B_{\perp}u=0$; alternately, we can choose any positive block $B$ in (48) and any value of $A_{12}$ satisfying (50). A naturally interesting question is then: for what fit parameters $\kappa_{k},\varepsilon_{k},\gamma_{k}$ will there be a choice of $u_{j},B_{\perp}$ such that we obtain $\Gamma>0$? We were unable to find a general answer to this question, but we explore it further for the simplest case of 2 pseudomodes in Appendix E. For that case, we show that it is possible to have $\Gamma>0$ if and only if the effective spectral density from the fit parameters is positive everywhere.
@@ III.3 Numerical Implementation
We choose this spectral density because fermionic system coupled to an infinite, homogenous wire will have a semi-elliptical spectral density Ferracin et al. (2024). We add the wide Lorentzian to represent how no physical system has true hard cutoff frequencies in its spectral density, and because the methods of section III.1 struggle to fit the sharp corners of a semicircular spectral density without producing an effective spectral density that is negative somewhere.
@@
As an example, we construct a 6-mode fit to the spectral density (56), using $\Delta=0.04$ and $\delta=6.0$. When optimizing over $u,B$, we obtain five positive rates and one negative rate. The spectral density and memory kernel and their corresponding fits are shown in figure 4 (a) and (b) respectively, along with the effective fits that we obtain if we set the one negative rate to zero. We compare our results with an optimized fit involving only uncoupled modes. While in this case we are unable to obtain a fit with only nonnegative rates, our modified fit where the negative rates are discarded still fits the true memory kernel quite well at both short and long times with only minor deviations at the peaks and troughs of $\chi(t)$ at intermediate times.
@@ IV.1 Proof of Nonconvergence
Figure 6: Effective spectral density for pseudomodes in a diagonal configuration (blue) and for modes in 2x2 non-diagonalizable blocks (pink), compared to the error expected in the infinite-mode limit at the pseudomode energies (above $J_{\rm eff}$) and halfway between them (below $J_{\rm eff}$, for two example spectral densities. 100 pseudomodes are used in each fit in each figure, creating an effective spectral density that is either the sum of 100 Lorentzians (100 uncoupled modes) or 50 squared Lorentzians (100 modes, coupled as 50 pairs), with $\omega_{m,M}=\pm 1.4$ for both figures. Note that where the effective spectral density differs from the infinite-mode prediction, it is where the second derivative of the true spectral density is large, so the first-order Taylor approximation (61) is locally a poor approximation. The left spectral density is semi-elliptical, and the right is a sum of two Lorentzians.
@@ IV.2 Many-Mode Approach with Non-diagonalizability
We can do slightly better in the large $n$ limit at the cost of doubling the number of modes by replacing the pseudomodes above with 2x2 non-diagonalizable blocks, with one of the two pseudomodes in each block coupled to the system and the other coupled to the residual bath as shown in Fig. 3, such that each block contributes a squared Lorentzian to the effective spectral density as in (36). The steps above are modified as:
@@
    Construct $\Lambda$ and $\Gamma$ as block-diagonal, with $n/2$ 2x2 blocks $\Lambda^{q}$ and $\Gamma^{q}$ of the form:
@@
There is still error in the large $n$ limit for this method, but it is smaller than that for the one-layer version. If we proceed analogously to before, we obtain at the pseudomode energies
@@
We can numerically compare the two methods to see these differences in error, as shown in Fig. 6. This oscillating error can be further mitigated by increasing $\gamma$ and adjusting the elements of $\zeta$ accordingly. However, while this reduces the magnitude of the oscillations, it makes the approximation (61) (and the corresponding approximation in the two-layer case) worse, and fails to fit the spectral density anywhere its second derivative is too large. It is likely that this issue could be mitigated by further refining this many-mode fitting procedure.
@@ V Connection to Scattering Theory
Here $\alpha,\beta\in\{L,R\}$, so this includes both transmission and reflection. The quantity $G$ is the single particle delayed non-equilibrium Green’s function (NEGF), which is an $n_{S}\times n_{S}$ matrix given by
@@
$J_{\alpha k}^{\rm res}(\omega)=2\pi\sum_{p\in(\alpha,k)}\tau_{\alpha k,p}^{2}\delta(\omega-\omega_{p})\,$, (77)
@@
$Q(\omega)=\begin{bmatrix}\Lambda_{L}+\Sigma_{L}(\omega)&0&\zeta_{L}\\[5.69046pt] 0&\Lambda_{R}+\Sigma_{R}(\omega)&\zeta_{R}\\[5.69046pt] \zeta_{L}^{\dagger}&\zeta_{R}^{\dagger}&H_{S}\end{bmatrix}\,$, (79)
@@
Using the fact that the matrices $\mathcal{X}_{\alpha k}$ in (80) are all sparse, we can write down four cases of interest:
@@
(78) means that $\mathcal{G}$ is the solution of the equations
@@
Next we express the four quantities $\mathcal{G}_{LL},\mathcal{G}_{RR},\mathcal{G}_{LR},\mathcal{G}_{RL}$ [which appear in (82)-(85)] in terms of $\mathcal{G}_{S}$. Combining (93) with (98) and (94) with (96) we get
@@ VI Conclusion
The method of pseudomodes is a useful approach to studying the dynamics of open quantum systems with strong coupling or structured environments. We have explored a number of subtleties that arise in the implementation of this technique, including the differences in effective spectral densities for coupled and uncoupled pseudomodes, and the appearance of new terms in the effective spectral density when the pseudomodes’ non-Hermitian Hamiltonian is non-diagonalizable. We have explored the pseudomode inversion problem and presented for the first time a method for finding pseudomode parameters that exactly match an effective fit to a spectral density. Our inversion method also sheds further light on the vast number of free parameters involved in the pseudomode method. We have also shown a connection between pseudomodes and scattering theory for non-interacting systems and shown that they produce the same effective spectral densities.

### 22 May 2026 v3

@@ I Introduction
The success of the pseudomodes approach, however, relies on the careful design of their geometry and parameters so as to properly mimic the real environment. Implementing this entails various subtleties, which are often overlooked. The goal of this paper is to address said subtleties. In section II, we consider the effective spectral density that arises from a particular configuration of pseudomodes. Depending on the non-Hermitian Hamiltonian generated by the pseudomodes’ Hamiltonian and the residual baths, three possible cases emerge. If none of the pseudomodes couple to each other, then one obtains an effective spectral density that is a sum of Lorentzians, while if the pseudomodes couple to each other there are additional terms in the effective spectral density which we call “Anti-Lorentzian”. We take this one step further, and show that one can obtain terms which are neither Lorentzian nor Anti-Lorentzian by using a non-diagonalizable non-Hermitian Hamiltonian for the pseudomodes, a case which to the best of our knowledge has not been previously characterized in detail.
@@ I.1 Model System
$-\sum_{\alpha}\sum_{j\in S}\intop_{0}^{t}dt^{\prime}\chi_{\alpha,ij}(t-t^{\prime})c_{j}(t^{\prime})$
@@
$C^{+}_{\alpha,ij}(t)=\intop\frac{d\omega}{2\pi}e^{i\omega t}J_{\alpha,ij}(\omega)f_{\alpha}(\omega)\,$, (6)
$C^{-}_{\alpha,ij}(t)=\intop\frac{d\omega}{2\pi}e^{-i\omega t}J_{\alpha,ij}(\omega)\big(1-f_{\alpha}(\omega)\big)\,$, (7)
@@
$\sum_{\alpha}\sum_{i\in S}c_{i}^{\dagger}B_{\alpha i}+B_{\alpha i}^{\dagger}c_{i}\,$, (8)
@@ I.2 Pseudomode Mapping
The first line of (I.2) describes the possible interactions between pseudomodes of the same bath, described by the $n_{\alpha}\times n_{\alpha}$ matrix $\Lambda_{\alpha}$ for each bath. If $\Lambda$ is diagonal then none of the pseudomodes couple to each other, and we have the simpler configuration shown in Fig. 1(b). The second line describes how each pseudomode interacts with the central system through a set of $n_{\alpha}\times n_{S}$ matrices $\zeta_{\alpha}$. Finally, the last line describes a set of infinite-dimensional residual reservoirs, one for each pseudomode $a_{\alpha k}$, labeled by the pair of indices $(\alpha,k)$ and composed of a set of modes $b_{\alpha k,p}$. The residual baths then have spectral densities given by
@@ II Computation of Effective Spectral Density
While all the results shown in this section follow straightforwardly from (14) and (16), we show them explicitly, as we will use these results for the remainder of the paper. Writing these results explicitly also makes clear the difference in the effective spectral densities that can be obtained using coupled modes as opposed to uncoupled modes, and why it is often desirable to use coupled modes. Additionally, we explore the appearance of new terms in the effective spectral density when the pseudomodes’ non-Hermitian Hamiltonian is non-diagonalizable.
@@ II.2 Diagonalizable Case
Using a diagonalizable but not diagonal $W$ often allows for a much better fit to structured spectral densities than using a diagonal $W$, as the negativities and anti-symmetry from the Anti-Lorentzian terms allow one to capture features much more easily. Figure 2 compares 6-mode fits to the semi-elliptical spectral density, showing how the diagonalizable version matches the true spectral density much more closely, particularly around the sharp corners at $\omega=\pm 1$. For a general memory kernel, the fitting error $\epsilon$ over a time window $T$ scales as $\mathrm{poly}(T/\epsilon)$ for uncoupled (diagonal) modes, but $\mathrm{polylog}(T/\epsilon)$ for diagonalizably coupled modes Huang et al. (2026); Thoenniss et al. (2025).
@@ II.3 Non-diagonalizable Case
Example: $\times2$ non-diagonalizable Case
Consider the case where we map bath $\alpha$, which couples to only one system site (so the spectral density matrix is $1\times 1$) to two pseudomodes, so $W_{\alpha}$ is $2\times 2$. If we want $W_{\alpha}$ to be non-diagonalizable, the most general form we can pick is:
@@ III Few-Mode Fitting
The most common approach to few-mode fitting is to first fit the memory kernel in terms of the parameters $\kappa_{\alpha k,ij},\varepsilon_{\alpha k},\gamma_{\alpha k}$, and then to try to choose pseudomode parameters $\Lambda,\Gamma,\zeta$ that give rise to the same effective memory kernel. We focus on this method. One might alternately choose to fit the correlation functions (6) (or their Fourier transforms), thus encoding the temperatures and chemical potentials of the original baths into the pseudomode parameters, rather than the residual baths Pleasance et al. (2020); Menczel et al. (2024). Few-mode fitting has also been performed by brute-force optimization over the pseudomode parameters without first fitting the memory kernel Feist et al. (2020); Medina et al. (2021); Sánchez-Barquilla et al. (2022); doing so guarantees a fit with physical pseudomode parameters (i.e, $\Gamma>0$), which we shall see is not always guaranteed otherwise, but is much less efficient computationally. A third method was recently proposed in Ref. Huang et al. (2026), which instead maps (16) onto the so-called control-theoretic realization problem Mayo and Antoulas (2007); Antoulas and Anderson (1986) and also guarantees a fit with physical pseudomode parameters.
@@ III.1 Fitting the Memory Kernel
We fit the memory kernel from each bath separately. We first fit each element of $\chi_{\alpha,ij}^{\rm eff}(t)$ as a sum of complex exponentials using techniques from discrete signal processing, such as Prony’s method Prony et al. (1795); Wilson et al. (2019); Almunif et al. (2020); Chen et al. (2022) or the ESPRIT algorithm Roy et al. (1986); Paulraj et al. (1986); Sahnoun et al. (2017) (see Appendix B). These methods work by sampling the true $\chi(t)$ at discrete, evenly spaced times, and then fitting those discrete points with complex exponentials. We found the best results by using ESPRIT, as it is more stable with respect to the set of points at which $\chi$ is sampled than Prony’s method. This gives us the set of $\kappa_{\alpha k,ij},\varepsilon_{\alpha k},\gamma_{\alpha k}$ from which to construct the effective memory kernel (23). Since the exponents in $\chi_{\alpha,ij}^{\rm eff}$ are independent of $i,j$, in the end each $i,j$ element of the memory kernel matrix should be fit by a sum of exponentials with the same exponents but different amplitudes $\kappa_{\alpha k,ij}$. Caution must be used here, since the fit obtained may produce a spectral density that is not positive everywhere or even has non-decaying exponentials. Note, also, that some spectral densities will be easier to fit than others. In particular, if we have a wideband $J(\omega)=\Gamma_{0}$ for $\omega_{m}\leq\omega\leq\omega_{M}$ and 0 otherwise, then the corresponding memory kernel is
@@ III.2 Obtaining Pseudomode Parameters
where $M$ is a diagonal matrix whose entries are the eigenvalues of $W$ obtained from fitting the memory kernel. While the eigenvalues of $W$ are specified, $S$ is not. Note that if we transform $W$ by some unitary transformation, then as long as we transform $\zeta$ by the same unitary we will obtain the same effective memory kernel (23):
@@
It is worth emphasizing again here the sheer amount of freedom involved in this process. In (43) we can choose anything for the $u_{j}$. Then in (44), $B_{\perp}$ can by any positive-definite $n\times n$ matrix satisfying $B_{\perp}u=0$; alternately, we can choose any positive block $B$ in (47) and any value of $A_{12}$ satisfying (49). A naturally interesting question is then as follows: For what fit parameters $\kappa_{k},\varepsilon_{k},\gamma_{k}$ will there be a choice of $u_{j},B_{\perp}$ such that we obtain $\Gamma>0$? We were unable to find a general answer to this question, but we explore it further for the simplest case of two pseudomodes in Appendix E. For that case, we show that it is possible to have $\Gamma>0$ if and only if the effective spectral density from the fit parameters is positive everywhere.
@@ III.3 Numerical Implementation
We choose this spectral density because fermionic system coupled to an infinite, homogeneous wire will have a semi-elliptical spectral density Ferracin et al. (2024). We add the wide Lorentzian to represent how no physical system has true hard cutoff frequencies in its spectral density, and because the methods of section III.1 struggle to fit the sharp corners of a semicircular spectral density without producing an effective spectral density that is negative somewhere.
@@
As an example, we construct a six-mode fit to the spectral density (55), using $\Delta=0.04$ and $\delta=6.0$. When optimizing over $u,B$, we obtain five positive rates and one negative rate. The spectral density and memory kernel and their corresponding fits are shown in Figs. 4(a) and 4(b) respectively, along with the effective fits that we obtain if we set the one negative rate to zero. We compare our results with an optimized fit involving only uncoupled modes. While in this case we are unable to obtain a fit with only nonnegative rates, our modified fit where the negative rates are discarded still fits the true memory kernel quite well at both short and long times with only minor deviations at the peaks and troughs of $\chi(t)$ at intermediate times.
@@ IV.1 Proof of Nonconvergence
Figure 6: Effective spectral density for pseudomodes in a diagonal configuration (blue) and for modes in $2\times 2$ non-diagonalizable blocks (pink), compared to the error expected in the infinite-mode limit at the pseudomode energies (above $J_{\rm eff}$) and halfway between them (below $J_{\rm eff}$), for two example spectral densities. 100 pseudomodes are used in each fit in each figure, creating an effective spectral density that is either the sum of 100 Lorentzians (100 uncoupled modes) or 50 squared Lorentzians (100 modes, coupled as 50 pairs), with $\omega_{m,M}=\pm 1.0$ for both figures. Note that where the effective spectral density differs from the infinite-mode prediction, it is where the second derivative of the true spectral density is large, so the first-order Taylor approximation (60) is locally a poor approximation. The left spectral density is semi-elliptical, and the right is a sum of two Lorentzians. The inset to (a) depicts the corresponding error in the time domain for the fits depicted in the main figure, as well as for a fit involving 50 uncoupled modes. The rapidly oscillating error in frequency space corresponds to a sharp peak at a large time in time space.
@@ IV.2 Many-Mode Approach with Non-diagonalizability
We can do slightly better in the large-$n$ limit at the cost of doubling the number of modes by replacing the pseudomodes above with $2\times 2$ non-diagonalizable blocks, with one of the two pseudomodes in each block coupled to the system and the other coupled to the residual bath as shown in Fig. 3, such that each block contributes a squared Lorentzian to the effective spectral density as in (35). The steps above are modified as:
@@
    Construct $\Lambda$ and $\Gamma$ as block-diagonal, with $n/2$ $2\times 2$ blocks $\Lambda^{q}$ and $\Gamma^{q}$ of the form:
@@
There is still error in the large-$n$ limit for this method, but it is smaller than that for the one-layer version. If we proceed analogously to before, we obtain at the pseudomode energies
@@
We can numerically compare the two methods to see these differences in error, as shown in Fig. 6. This rapidly oscillating error in the frequency domain becomes, in the time domain, a large sharply-peaked error around some time $T$ which increases with the number of pseudomodes used, as shown in the inset to Fig. 6(a). This oscillating error can be further mitigated by increasing $\gamma$ and adjusting the elements of $\zeta$ accordingly. However, while this reduces the magnitude of the oscillations, it makes the approximation (60) (and the corresponding approximation in the two-layer case) worse, and fails to fit the spectral density anywhere its second derivative is too large. It is likely that this issue could be mitigated by further refining this many-mode fitting procedure.
@@ V Connection to Scattering Theory
Here $\alpha,\beta\in\{L,R\}$, so this includes both transmission and reflection. The quantity $G$ is the single-particle delayed non-equilibrium Green’s function (NEGF), which is an $n_{S}\times n_{S}$ matrix given by
@@
$Q(\omega)=\begin{bmatrix}\Lambda_{L}+\Sigma_{L}(\omega)&0&\zeta_{L}\\ 0&\Lambda_{R}+\Sigma_{R}(\omega)&\zeta_{R}\\ \zeta_{L}^{\dagger}&\zeta_{R}^{\dagger}&H_{S}\end{bmatrix}\,$, (78)
@@
Using the fact that the matrices $\mathcal{X}_{\alpha k}$ in (79) are all sparse, we can write four cases of interest:
@@
Equation (77) means that $\mathcal{G}$ is the solution of the equations
@@
Next we express the four quantities $\mathcal{G}_{LL},\mathcal{G}_{RR},\mathcal{G}_{LR},\mathcal{G}_{RL}$ [which appear in (81)–(84)] in terms of $\mathcal{G}_{S}$. Combining (92) with (97) and (93) with (95) we get
@@ VI Conclusion
The method of pseudomodes is a useful approach to studying the dynamics of open quantum systems with strong coupling or structured environments. We have explored a number of subtleties that arise in the implementation of this technique, including the differences in effective spectral densities for coupled and uncoupled pseudomodes and the appearance of new terms in the effective spectral density when the pseudomodes’ non-Hermitian Hamiltonian is non-diagonalizable. We have explored the pseudomode inversion problem and presented a method for finding pseudomode parameters that exactly match an effective fit to a spectral density. Our inversion method also sheds further light on the vast number of free parameters involved in the pseudomode method. We have also shown a connection between pseudomodes and scattering theory for non-interacting systems and shown that they produce the same effective spectral densities.